OpenStudy (anonymous):

Could you possibly help me figure out how to find the length of each arc. Answer to three decimal places. Arc JK:m arc JK=79°, r=56 cm Please help.

OpenStudy (anonymous):



OpenStudy (kittiwitti1):

\[s=r\theta\]s being arc length, r being radius and theta being the angle in RADIANS (so you have to convert)

OpenStudy (kittiwitti1):

btw this is how you convert:\[79°\times\frac{\pi}{180°}\]

OpenStudy (anonymous):

Thank you so much that is very helpful

OpenStudy (kittiwitti1):

No problem ☻